If P(A) = 0.25, P(B) = 0.35, and P(A intersection B) = 0.20 then, P(A union B) =
Answer #1

P(A union B) = P(A) + P(B) - P(A intersection B)

Substitute given values

=>P(A union B) = P(A) + P(B) - P(A intersection B)

=0.25+0.35 - 0.20

=0.40